OpenStudy (anonymous):

WILL GET MEDAL! find the solutions to the systems of nonlinear equations. 1. y=2x x^2+y^2=15 2. y=3x x^2+y^2=40

OpenStudy (anonymous):

halp pls

OpenStudy (anonymous):

Substitution.

OpenStudy (anonymous):

\[ x^2+y^2\Bigg|_{ y=2x}=x^2+(2x)^2 = 5x^2=15 \]Solve for \(x\) here, it should be simple.

OpenStudy (anonymous):

x=(sqrt)3

OpenStudy (anonymous):

y=2(sqrt 3)?

OpenStudy (anonymous):

\(x=\pm \sqrt 3\)

OpenStudy (anonymous):

Would I give the answer for both, x and y?

OpenStudy (anonymous):

Yes, so this one has two solutions.

OpenStudy (anonymous):

Alright, thanks so much for your help and patience!

OpenStudy (anonymous):

Now, for the second equation, would these be the answers? x=2 y=6

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

Yeah, again there is a positive and negative solution.

OpenStudy (anonymous):

Thanks!